4x+5y=140
3x-5y=105

Answer by phillywily(18)

4x+5y=140
+)3x-5y=105
7x=245
x=35
5y=140-4x
5y=140-4(35)
5y=140-140
5y=0
y=0
To check:
4x+5y=140
4(35)+5(0)=140
140+0=140
,
3x-5y=105
3(35)-5(0)=105
105-0=105
Hence, x=35 and y=0.